# Diameter Bounds for Friends-and-Strangers Graphs

### v1

@@ 5. General Diameter Bounds Depending on Minimum Degree
For the first fact, we cite another theorem in Bangachev, which says that $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected if $\delta(Y)>\frac{n}{2}$. Hence, by Theorem˜1.3 we conclude that the diameter of $\mathsf{FS}(\mathsf{Star}_{|Q|},Y|_{Q}))$ is at most $O(|Q|^{4})\leq O(n^{4})$. Furthermore, $X|_{\sigma(Q)}$ clearly has a subgraph isomorphic to $\mathsf{Star}_{|Q|}$, because $u$ is adjacent to every vertex in this graph. Hence, we conclude that $\mathsf{FS}(\mathsf{Star}_{|Q|},Y|_{Q}))$ is a subgraph of $\mathsf{FS}(X|_{\sigma(Q)},Y|_{Q})$, and the latter graph has diameter $O(n^{4})$. The assertion follows, because this graph is connected and we can therefore clearly switch the positions of $\hat{u}$ and $\hat{v}$ in $O(n^{4})$ moves.
@@
We remark that Bangachev [3] proves another connectivity result of a similar flavor as the connectivity analog of Theorem˜1.5; see Figure˜1 for a visual comparison of the regimes that these results apply to. We state this result below for completeness.
@@
[3, Theorem 1.4] Let $X$ and $Y$ be graphs on $n$ vertices. If $\delta(X)>n/2$, $\delta(Y)>n/2$, and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$, then $\mathsf{FS}(X,Y)$ is connected.
@@ 6. Diameter Bounds for Erdős-Rényi Random Graphs
In this section we prove Theorem˜1.4, which we restate below for convenience. See 1.4
As in the work of Alon, Defant and Kravitz [2], our proof relies on the notion of Wilsonian graphs, as defined below. Theorem˜4.3 then shows that if $Y$ is Wilsonian, then $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected. We will use this fact in our argument.
@@
We’ll split this into two cases:
@@
    We first consider the case $\left\lvert M\right\rvert\leq 20$. Fix any nonempty subset $M\subseteq V(X)$ of cardinality at most 20. Then for any vertex in $M$, the number of edges it has with the $n-\left\lvert M\right\rvert\geq n-20$ vertices in $V(X)\setminus M$ is at most 1, namely an edge to $v$. We’ll use Chernoff bounds to bound the probability that any fixed $x\in M$ has at most 1 edge to $V(X)\setminus M$. Note that these events for different $x\in M$ are mutually independent.
@@
    $\mu=20\log n-\frac{20M\log n}{n}\geq 20\log n-\frac{400\log n}{n},$
    so if we define $\delta=\frac{\mu-1}{\mu}$, then the probability $x\in M$ has at most 1 edge to $V(X)\setminus M$ is at most $e^{-\delta^{2}\mu/2}$. For large enough $n$, we have $\delta^{2}\geq 0.991$, so this value is bounded from above by
    $e^{-0.991\mu/2}<n^{-9.9},$
    where this inequality holds for all sufficiently large $n$.
@@
    $\sum_{r=21}^{n/2-1}\binom{n}{r}\cdot(n-r)\cdot(1-p)^{r(n-1-r)},$
    where the first term represents the number of ways to choose the subgraph $M$, the second term represents the number of ways to choose $v$, and the third term represents the probability that none of the edges between $M$ and $N$ exist. We can bound this from above by
@@
We first note that, by Lemma˜6.2, $X$ is connected with probability $1-o(n^{-8})$. Moreover, with high probability the degree of every vertex in $X$ is in $(1-\epsilon)m,(1+\epsilon)m]$. In particular, for a fixed vertex $v\in V(X)$, the expected degree of $v$ in $X$ is $m$, and therefore using Chernoff bounds, the probability that $v$’s degree falls outside of the range $(1-\epsilon)m,(1+\epsilon)m]$ is at most $\operatorname{exp}(-\epsilon^{2}m/2)=n^{-5}$. Taking the union bound over the $n$ vertices, we get that with probability at least $1-n^{-4}$, every vertex in $X$ has degree in $(1-\epsilon)m,(1+\epsilon)m]$ and $X$ is connected. We will condition on this event for the rest of the argument.
@@
    Then, by Lemma˜6.2, it follows that $H_{Y}$ is not biconnected with probability at most $o(t^{-8})=o(n^{-8/2})$, where we used $t\geq(1-\epsilon)(n-1)p=\Omega(\sqrt{n\log n})$. Therefore, $H_{Y}$ is biconnected with probability $1-o(n^{-4})$.
@@
    $\sum_{j=0}^{t}{t\choose j}(1-q)^{{j\choose 2}+{{t-j}\choose 2}}\leq 2^{t}(1-q)^{t^{2}/4}<2^{t}e^{-qt^{2}/4}<\operatorname{exp}(t-t\log n)=o(n^{-4}),$
    where we used the fact that $qt\geq(1-\epsilon)(n-1)pq>4\log n$ and $t>5$.
@@
Taking the union bound over the three bad events and applying Definition˜6.1, we get that the probability that $H_{Y}$ is not Wilsonian is at most $o(n^{-4})$, and so we are done. ∎
@@ 7. Open Questions
Jeong’s friends-and-strangers graph construction for the equality case in the above question has many connected components. This raises a natural followup question.
Question 7.2 ([7]).
@@
We’re also interested in improving our diameter bound on $\mathsf{FS}(\mathsf{Star}_{n},Y)$. In this paper, we showed that the bound is $O(n^{4})$, but we have only been able to show that it is bounded from below by $\Omega(n^{3})$. This raises the following natural question.
Question 7.3.
Do there exist graphs $Y$ for which $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has diameter $\Omega(n^{4})$? If not, what is the maximum possible diameter $\mathsf{FS}(\mathsf{Star}_{n},Y)$ can have?
We can ask a similar but more general question about trees, as follows.
Question 7.4.
@@
We would also like to extend our results in Section˜5, as shown in Figure˜1. We showed that the diameter of $\mathsf{FS}(X,Y)$ is polynomial in $n$ when $\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n$. On the other hand, Bangachev [3] showed that $\mathsf{FS}(X,Y)$ is connected when $\delta(X),\delta(Y)>\frac{n}{2}$ and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$. This result covers many graphs which are not shown to have polynomial diameter in Theorem˜1.5. We ask whether a counterpart of Theorem˜1.5 extends to this regime.
Question 7.5.
@@
In Section˜6, we showed that when $X$ and $Y$ are Erdős-Rényi random graphs satisfying certain conditions, the distance between any two permutations in $\mathsf{FS}(X,Y)$ is $O(n^{6})$ with high probability. We’re curious to know if in the same regime, one can establish polynomial diameter bounds with high probability.
Question 7.7.

### v2

@@ 4. General Diameter Bounds Depending on Minimum Degree
For the first fact, we cite another theorem in Bangachev, which says that $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected if $\delta(Y)>\frac{n}{2}$. Hence, by Theorem˜2.2 we conclude that the diameter of $\mathsf{FS}(\mathsf{Star}_{|Q|},Y|_{Q})$ is at most $O(|Q|^{3})\leq O(n^{3})$. Furthermore, $X|_{\sigma(Q)}$ clearly has a subgraph isomorphic to $\mathsf{Star}_{|Q|}$, because $u$ is adjacent to every vertex in this graph. Hence, we conclude that $\mathsf{FS}(\mathsf{Star}_{|Q|},Y|_{Q})$ is a subgraph of $\mathsf{FS}(X|_{\sigma(Q)},Y|_{Q})$, and the latter graph has diameter $O(n^{3})$. The assertion follows, because this graph is connected and we can therefore clearly switch the positions of $\hat{u}$ and $\hat{v}$ in $O(n^{3})$ moves.
@@
We remark that Bangachev [2] proves another connectivity result of a similar flavor as the connectivity analog of Theorem˜1.3; see Figure˜1 for a visual comparison of the regimes that these results apply to. We state this result below for completeness.
@@
[2, Theorem 1.4] Let $X$ and $Y$ be graphs on $n$ vertices. If $\delta(X)>n/2$, $\delta(Y)>n/2$, and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$, then $\mathsf{FS}(X,Y)$ is connected.
@@ 5. Diameter Bounds for Erdős-Rényi Random Graphs
In this section we prove Theorem˜1.5, which we restate below for convenience. See 1.5
As in the work of Alon, Defant and Kravitz [1], our proof relies on the notion of Wilsonian graphs, as defined below. Theorem˜2.3 then shows that if $Y$ is Wilsonian, then $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected. We will use this fact in our argument.
@@
We will split this into two cases:
@@
    We first consider the case $\left\lvert M\right\rvert\leq 20$. Fix any nonempty subset $M\subseteq V(X)$ of cardinality at most 20. Then for any vertex in $M$, the number of edges it has with the $n-\left\lvert M\right\rvert\geq n-20$ vertices in $V(X)\setminus M$ is at most 1, namely an edge to $v$. We will use Chernoff bounds to bound the probability that any fixed $x\in M$ has at most 1 edge to $V(X)\setminus M$. Note that these events for different $x\in M$ are mutually independent.
@@
    $\mu=20\log n-\frac{20M\log n}{n}\geq 20\log n-\frac{400\log n}{n}.$
    Therefore, if we define $\delta=\frac{\mu-1}{\mu}$, then the probability $x\in M$ has at most 1 edge to $V(X)\setminus M$ is at most $e^{-\delta^{2}\mu/2}$. For large enough $n$, we have $\delta^{2}\geq 0.991$, so this value is bounded from above by
    $e^{-0.991\mu/2}<n^{-9.9}.$
@@
    $\sum_{r=21}^{n/2-1}\binom{n}{r}\cdot(n-r)\cdot(1-p)^{r(n-1-r)}.$
    Here, the first term represents the number of ways to choose the subgraph $M$, the second term represents the number of ways to choose $v$, and the third term represents the probability that none of the edges between $M$ and $N$ exist. We can bound this from above by
@@
We first note that, by Lemma˜5.2, $X$ is connected with probability $1-o(n^{-8})$. Moreover, with high probability the degree of every vertex in $X$ is in $(1-\epsilon)m,(1+\epsilon)m]$. In particular, for a fixed vertex $v\in V(X)$, the expected degree of $v$ in $X$ is $m$, and therefore using Chernoff bounds, the probability that $v$’s degree falls outside of the range $(1-\epsilon)m,(1+\epsilon)m]$ is at most $\operatorname{exp}(-\epsilon^{2}m/2)=n^{-5}$. Taking the union bound over the $n$ vertices, we get that with probability at least $1-n^{-4}$, every vertex in $X$ has degree in $(1-\epsilon)m,(1+\epsilon)m]$ and $X$ is connected. We will condition on this event for the rest of the argument.
@@
    Then, by Lemma˜5.2, it follows that $H_{Y}$ is not biconnected with probability at most $o(t^{-8})=o(n^{-8/2})$, where we used $t\geq(1-\epsilon)(n-1)p=\Omega(\sqrt{n\log n})$. Therefore, $H_{Y}$ is biconnected with probability $1-o(n^{-4})$.
@@
    $\sum_{j=0}^{t}\binom{t}{j}(1-q)^{\binom{j}{2}+\binom{t-j}{2}}\leq 2^{t}(1-q)^{t^{2}/4}<2^{t}e^{-qt^{2}/4}<\operatorname{exp}(t-t\log n)=o(n^{-4}).$
    Note that we used the fact that $qt\geq(1-\epsilon)(n-1)pq>4\log n$ and $t>5$.
@@
Taking the union bound over the three bad events and applying Definition˜5.1, we get that the probability that $H_{Y}$ is not Wilsonian is at most $o(n^{-4})$, and so we are done. ∎
@@ 6. Open Questions
Jeong’s friends-and-strangers graph construction for the equality case in the above question has many connected components. This raises a natural follow-up question.
Question 6.2 ([6]).
@@
[7] shows that the diameter of the connected components of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is bounded polynomially. We can ask a similar but more general question about trees, as follows.
Question 6.3.
@@
We would also like to extend our results in Section˜4, as shown in Figure˜1. We showed that the diameter of $\mathsf{FS}(X,Y)$ is polynomial in $n$ when $\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n$. On the other hand, Bangachev [2] showed that $\mathsf{FS}(X,Y)$ is connected when $\delta(X),\delta(Y)>\frac{n}{2}$ and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$. This result covers many graphs which are not shown to have polynomial diameter in Theorem˜1.3. We ask whether a counterpart of Theorem˜1.3 extends to this regime.
Question 6.4.
@@
In Section˜5, we showed that when $X$ and $Y$ are Erdős-Rényi random graphs satisfying certain conditions, the distance between any two permutations in $\mathsf{FS}(X,Y)$ is $O(n^{6})$ with high probability. We are curious to know if in the same regime, one can establish polynomial diameter bounds with high probability.
Question 6.6.